# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

### v1

@@ 2.3.1 Spectroscopy from Subaru/IRD
We performed high-resolution spectroscopy for TOI-6008 with the InfraRed Doppler (IRD) instrument (Tamura et al. 2012; Kotani et al. 2018) mounted at the Subaru 8.2m telescope. IRD can obtain a near-infrared spectrum between from 950 nm to 1740 nm at one exposure with a spectral resolution of $R\approx 70,000$. Between UT 2023 July 31 and August 6, a total of 15 frames were collected for TOI-6008, for which integration times were set to 1200–1500 seconds. We also injected the laser-frequency comb (LFC) into a secondary fiber so that we can trace the temporal variation of the instrumental profile.
IRD data were reduced following the reduction procedure in Hirano et al. (2020), where custom reduction codes as well as IRAF (Tody 1993) were used. For each frame, one-dimensional stellar and LFC spectra were separately extracted by this reduction. The reduced stellar spectra typically have signal-to-noise (S/N) ratios of 60–70 per pixel at 1000 nm. Those spectra were then subjected to the radial-velocity (RV) measurements with the standard RV pipeline for IRD (Hirano et al. 2020), and we extracted relative RVs for individual frames using the template-matching technique. The resulting RV measurements and curve are presented in Table 2 and Figure 6, respectively; the RV internal errors returned by the pipeline were typically $\approx 3$ m s-1.
@@
Table 2: Radial-velocities measurements for TOI-6008 obtained from Subar-8.2m/IRD spectrograph.
Figure 6: RVs obtained with the Subaru-8.2m/IRD spectrograph for TOI-6008. a, RVs observations versus time. b, RVs observations phase-folded. The best-fitting model is superimposed in red.
@@ 2.3.2 Spectroscopy from TRES
Two reconnaissance spectra of TOI-5720 were obtained on UT December 19 and 24, 2022 using the Tillinghast Reflector Echelle Spectrograph (TRES; Fűrész (2008)) mounted on the 1.5m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory (FLWO) in Arizona. TRES is a fiber-fed spectrograph that observes in the 390–910 nm range with a resolving power of R=44,000. Spectra were extracted using methods described in Buchhave et al. (2010).
For the two TRES observations of TOI-5720, we measure radial velocities using a pipeline optimized for mid-to-late M dwarfs (see e.g. Pass et al. (2023) for further details). In brief, this pipeline performs a cross correlation between the observed spectra and an empirical mid-to-late M dwarf template using six TRES echelle orders in the range of 6400–7850 Å. The pipeline also produces carefully calibrated radial velocity uncertainties, taking into account the signal-to-noise ratio of the observed spectra, template mismatch, rotational broadening as well as the long-term stability of the spectrograph and errors in the barycentric correction. The two resulting RVs measurements are $35.692\pm 0.031$ km/s on UT December 19 and $35.684\pm 0.024$ km/s for UT December 24, and an upper limit on the projected rotational velocity is found to be $v\sin(i)<3.4$ km/s.
2.4 High-Resolution imaging
We assess the possibility of a blend of nearby stellar companions as part of our standard process for validating transiting exoplanets, that can dilute the transit light curves and underestimated planetary radii measurements (Ciardi et al. 2015). We observed TOI-5720, TOI-6008 and TOI-6086 with WIYN 3.5 m/NESSI, Palomar-5m0/PHARO and Robo-AO high-resolution imaging. Despite the combination of Gaia catalog and high-resolution imaging, there is no indication of any additional close stellar companions.
2.4.1 Palomar-5m0/PHARO
We observed TOI-5720, TOI-6008 and TOI-6086 with the PHARO instrument (Hayward et al. 2001) behind the natural guide star AO system P3K (Dekany et al. 2013) on UT 2023 June 07 and 08 in the narrow-band $Br-\gamma$ filter ($\lambda_{0}=2.1686;\Delta\lambda=0.0326$ $\mu$m). PHARO equipped with a 1024$\times$1024 HgCdTe HAWAII detector, with a pixel scale of 0.025 arcsec, resulting a total FOV of 25 arcsec. Each dither position was observed three times for a total of 15 frames; with an exposure time of 9.9s for TOI-5720, 2.8s for TOI-6008 and 9.9s for TOI-6086. The total on-source time was 148.5s for TOI-5720, 42.5s for TOI-6008 and 148.5s for TOI-6086. The science data were calibrated by sky-subtracted and flat-fielded. Further details of observations from the Palomar-5m0/PHARO are available in Hayward et al. (2001). No companion sources were detected for TOI-5720, TOI-6008 and TOI-6086 in the PHARO data (see Figure 7).
@@ 3 Stellar characterisation
Table 3: Spectroscopy, astrometry and photometry stellar properties of the host stars TOI-5720, TOI-6008 and TOI-6086. (1) Gaia EDR3; Gaia Collaboration et al. (2021a); (2) TESS Input Catalog; Stassun et al. (2018b); (3) UCAC4; Zacharias et al. (2012); (4) 2MASS; Skrutskie et al. (2006); (5) WISE; Cutri et al. (2021).
@@
Table 4: Planetary physical parameters determined from our global MCMC analysis (medians and $1\sigma$ error bar) for TOI-5720 b, TOI-6008 b and TOI-6086 b. a The equilibrium temperature derived assuming a null albedo. b Planetary mass and RV semi-amplitude values derived from RVs measurements collected by Subaru/IRD and TrES spectrographs for TOI-6008 b and TOI-5720 b, respectively. c Predicted mass and RV semi-amplitude values calculated using the mass-radius relationship from Chen & Kipping (2017). d TSM (Transmission spectroscopy metric) and ESM (emission spectroscopy metric) values calculated using equations 1 and 4 from Kempton et al. (2018).
@@
Table 5: Derived stellar parameters ($\rho_{\star}$, $M_{\star}$, $R_{\star}$, $L_{\star}$, $T_{\rm eff}$) and Limb-darkening coefficients for TOI-5720, TOI-6008 and TOI-6086 based on our global analysis.
@@ 3.2 IRTF/SpeX spectroscopic analysis
3.3 SED analysis and evolutionary models
We conducted an independent analysis of the broadband SED (spectral energy distribution) of the stars together with the Gaia DR3 parallax (with no systematic offset applied; see, e.g., Stassun & Torres 2021), to derive an empirical measurement of the stellar radii, following the same method described in Stassun & Torres (2016); Stassun et al. (2017, 2018a). We pulled the 2MASS $JHK_{S}$ magnitudes, the WISE W1–W4 magnitudes, the Pan-STARRS $grizy$ magnitudes, as well as the Gaia $G_{\rm BP}G_{\rm RP}$ magnitudes. Together, the available photometry spans the full stellar SED over the wavelength range 0.4–20 $\mu$m (see Figure 8).
Figure 8: Spectral Energy Distribution (SED) fit of TOI-5720 (left), TOI-6008 (middle) and TOI-6086 (right). The gray curves are the best-fitting PHOENIX atmosphere model, colored symbols with error-bars are the observed fluxes (horizontal bars represent the effective bandpasses), and black symbols are the model fluxes.
We performed a fit using PHOENIX stellar atmosphere models (Husser et al. 2013a), with the free parameters being the metallicity ([Fe/H]) and effective temperature ($T_{\rm eff}$) we set null extinction due to the close proximity of the stars to Earth. The resulting fits (Figure 8) have a best-fit $T_{\rm eff}=3325\pm 75$ K, $T_{\rm eff}=3075\pm 75$ K, and $T_{\rm eff}=3200\pm 75$ K for TOI-5720, TOI-6008, and TOI-6086 respectively, and a best-fit [Fe/H] = $0.0\pm 0.3$, $+0.3\pm 0.3$, and $-0.2\pm 0.3$ respectively, with a reduced $\chi^{2}$ of 1.8, 1.9, and 2.4 respectively (the latter star exhibits some UV excess that may indicate chromospheric activity). Integrating the model SED gives the bolometric flux at Earth, $F_{\rm bol}=4.027\pm 0.094\times 10^{-10}$ erg s-1 cm-2, $F_{\rm bol}=2.839\pm 0.101\times 10^{-10}$ erg s-1 cm-2, and $F_{\rm bol}=2.051\pm 0.072\times 10^{-10}$ erg s-1 cm-2 respectively. Taking the $F_{\rm bol}$ together with the Gaia parallax, gives the bolometric luminosity directly as $L_{\rm bol}=0.01615\pm 0.00019$ L, $L_{\rm bol}=0.004691\pm 0.000082$ L, and $L_{\rm bol}=0.00634\pm 0.00011$ L respectively. The stellar radii then follow from the Stefan-Boltzmann relation as $R_{\star}=0.383\pm 0.018$ R, $R_{\star}=0.242\pm 0.013$ R, and $R_{\star}=0.259\pm 0.013$ R respectively. Finally, we can estimate the stellar mass from the empirical mass-luminosity relation of Mann et al. (2019), giving $M_{\star}=0.383\pm 0.019$ M, $M_{\star}=0.230\pm 0.011$ M, and $M_{\star}=0.254\pm 0.013$ M respectively.
@@ 4.1 Data validation report for TESS
As we discribed in Section 2.1, the SPOC pipeline extracted the photometric measurements for TOI-5720, TOI-6008 and TOI-6086 using 2-minutes cadence, yielding orbital periods of $P=1.43446$ days, $P=0.85737$ day, and $P=1.38886$ days, and signal-to-noise ratios of $S/N=7.7$, $S/N=14.9$ and $S/N=8.1$, respectively. The TOI vetting team reviewed the SPOC Data Validation reports (Jenkins 2002; Jenkins et al. 2010, 2020) on May 26 2022, Sept 16 2022 and Sept 15 2022 (Guerrero et al. 2021) for TOI-5720, TOI-6008 and TOI-6086, respectively. The transit depths were found to be $1.03\pm 0.16$ ppt (for TOI-5720.01), $2.05\pm 0.27$ ppt (for TOI-6008.01) and $1.37\pm 0.24$ ppt (for TOI-6086.01), corresponding to a planet radius of $R_{p}=1.4\pm 0.3\pm$ R, $R_{p}=1.2\pm 0.2\pm$ R and $1.0\pm 1.7\pm$ R for TOI-5720.01, TOI-6008.01 and TOI-6086.01, respectively. All additional validation tests, including bootstrap, centroid offset, ghost, difference imaging centroid tests, as well as a search for discrepancies between odd and even transit depths were successfully passed. These conclusive tests prompted us to conduct further high-precision photometric follow-up and analysis to validate the planet candidates.
@@ 4.2 Vetting
While the transit method is one of the most fruitful technique for discovering exoplanets, several scenarios other than a transiting exoplanet can produce periodic dips in the brightness of a star. Numerous false positive scenarios exist in astronomical observations: i) stellar binary systems, where two stars sharing a common center might create a signal resembling a planetary transit if their sizes differ or if their eclipse is partial; ii) blended binary stars, wherein the target blends the light of an eclipsing binary obscuring the secondary eclipse and confusing the primary with a planetary transit; iii) other misleading signals from astrophysical sources (stellar spots, pulsations, rotation) or instrument-related factors (jitter noise, momentum dumps).
@@
Thus, the first essential step towards the confirmation of a transiting exoplanetary signal is the vetting. In addition to the SPOC DV reports for TOI 5720.01, TOI 6008.01 and TOI 6086.01, we conduct a uniform vetting analysis through DAVE (Discovery and Vetting of Exoplanets, Kostov et al. 2019) which allows to examine a transit event on a double-level analysis: the pixel level by a photocenter analysis and the light curve level by a flux time-series analysis. The centroid module generates a difference image by subtracting the overall in-transit image from the out-of-transit image. It determines the photocenter by fitting the TESS pixel response function to the image, performed for each detected transit. The average centroid position across events is calculated, along with its statistical significance, aiding the vetting process. An offset in the centroid position from the expected target star’s position flags the event as a False Positive (FP). However, challenging interpretation due to artifacts or low SNR in DAVE-produced difference images makes the photocenter analysis unreliable. In such cases we classify the signal as a planet candidate. The Modelshift module produces a phase-folded light curve with the best-fit trapezoid transit model. Its primary goal is to determine if the signal source is an eclipsing binary system. It showcases key transit features, including the average primary signal and other significant features. Additionally, it compares odd and even transits, measuring their statistical difference. It also evaluates the transit shape, offering a comprehensive assessment of the observed transit properties.
@@
In the vetting analysis of planetary transit candidates, reliability can be compromised when the SNR of the event is particularly low. This occurs because low SNR signals may be indistinguishable from noise or other sources of variability in the data, making it challenging to accurately identify genuine transit events. As a result, vetting efforts may be less effective in confirming or ruling out potential candidates in such cases. TOI 5720.01, TOI 6008.01 and TOI 6086.01 are three low SNR planet candidates with the first and the third being near the detection threshold (SNR=$7.1$, Borucki et al. 2011). However their transit depths on the phase-folded light curve are slightly above the noise level. Moreover, we did not flag any particular anomalies in the shape of the transit as well as no statistical significant evidence arose between the odd and the even transits. As the centroid module is concerned, TOI 5720.01 passed the photocenter test without any concern. The same test was inconclusive for both TOI 6008.01 and TOI 6086.01 since the TESS aperture masks used to extract their light curves are crowded with contaminating stars, resulting in an unreliable overall centroid measurement. However, for a few single clean centroid measurements the light photocenters are consistent with the position of both TOI 6008 and TOI 6086, as we will discuss in Sect.4.3.

### v2

@@ 2.3.1 Spectroscopy from Subaru/IRD
We performed high-resolution spectroscopy for TOI-6008 with the InfraRed Doppler (IRD) instrument (Tamura et al. 2012; Kotani et al. 2018) mounted at the Subaru 8.2m telescope. IRD can obtain a near-infrared spectrum between from 950 nm to 1740 nm at one exposure with a spectral resolution of $R\approx 70,000$. Between UT 2023 July 31 and August 6, a total of 15 frames were collected for TOI-6008, for which integration times were set to 1200–1500 seconds. We also injected the laser-frequency comb (LFC) into a secondary fiber so that we were able to trace the temporal variation of the instrumental profile.
IRD data were reduced following the reduction procedure in Hirano et al. (2020), where custom reduction codes as well as IRAF (Tody 1993) were used. For each frame, one-dimensional stellar and LFC spectra were separately extracted by this reduction. The reduced stellar spectra typically have (S/N) of 60–70 per pixel at 1000 nm. These spectra were then subjected to the radial velocity (RV) measurements with the standard RV pipeline for IRD (Hirano et al. 2020), and we extracted relative RVs for individual frames using the template-matching technique. The resulting RV measurements and curve are presented in Table 2 and Figure 6, respectively; the RV internal errors returned by the pipeline were typically $\sim 3$ m s-1.
@@
Table 2: Radial velocity measurements for TOI-6008 obtained from the Subar-8.2m/IRD spectrograph.
Figure 6: Radial velocities obtained with the Subaru-8.2m/IRD spectrograph for TOI-6008. a, RVs observations vs. time. b, Phase-folded RV observations. The best-fitting model is superimposed in red.
@@ 2.3.2 Spectroscopy from TRES
Two reconnaissance spectra of TOI-5720 were obtained on UT December 19 and 24, 2022 using the Tillinghast reflector Echelle Spectrograph (TRES; Fűrész (2008)) mounted on the 1.5m Tillinghast Reflector telescope at the Fred Lawrence Whipple Observatory (FLWO) in Arizona. TRES is a fiber-fed spectrograph that observes in the 390–910 nm range with a resolving power of R=44,000. Spectra were extracted using the methods described in Buchhave et al. (2010).
For the two TRES observations of TOI-5720, we measured the radial velocities using a pipeline optimized for mid- to late-M dwarfs (see, e.g., Pass et al. (2023) for further details). In brief, this pipeline performs a cross correlation between the observed spectra and an empirical mid- to late-M dwarf template using six TRES echelle orders in the range of 6400–7850 Å. The pipeline also produces carefully calibrated radial velocity uncertainties, taking into account the signal-to-noise ratio of the observed spectra, template mismatch, rotational broadening, and the long-term stability of the spectrograph and errors in the barycentric correction. The two resulting RV measurements are $35.692\pm 0.031$ km/s on UT December 19 and $35.684\pm 0.024$ km/s for UT December 24, and the upper limit on the projected rotational velocity is found to be $v\sin(i)<3.4$ km/s.
2.4 High-resolution imaging using Palomar-5m0/PHARO
We assessed the possibility of a blend of nearby stellar companions as part of our standard process for validating transiting exoplanets, which can dilute the transit light curves and underestimated planetary radii measurements (Ciardi et al. 2015). We observed TOI-5720, TOI-6008, and TOI-6086 with WIYN 3.5 m/NESSI, Palomar-5m0/PHARO, and Robo-AO high-resolution imaging. Despite the combination of Gaia catalog and high-resolution imaging, there is no indication of any additional close stellar companions.
We observed TOI-5720, TOI-6008, and TOI-6086 with the PHARO instrument (Hayward et al. 2001) behind the natural guide star AO system P3K (Dekany et al. 2013) on UT 2023 June 7 and 8 in the narrow-band $Br-\gamma$ filter ($\lambda_{0}=2.1686;\Delta\lambda=0.0326$ $\mu$m). PHARO is equipped with a 1024$\times$1024 HgCdTe HAWAII detector, with a pixel scale of 0.025 arcsec, resulting in a total FOV of 25 arcsec. Each dither position was observed three times for a total of 15 frames with an exposure time of 9.9s for TOI-5720, 2.8s for TOI-6008, and 9.9s for TOI-6086. The total on-source time was 148.5s for TOI-5720, 42.5s for TOI-6008, and 148.5s for TOI-6086. The science data were calibrated by sky-subtraction and flat-fielding. Further details of the observations from Palomar-5m0/PHARO are available in Hayward et al. (2001). No companion sources were detected for TOI-5720, TOI-6008, and TOI-6086 in the PHARO data (see Figure 7).
@@ 3 Stellar characterization
Table 3: Spectroscopy, astrometry and photometry stellar properties of the host stars TOI-5720, TOI-6008, and TOI-6086. (1) Gaia EDR3; Gaia Collaboration et al. (2021a); (2) TESS Input Catalog; Stassun et al. (2018b); (3) UCAC4; Zacharias et al. (2012); (4) 2MASS; Skrutskie et al. (2006); (5) WISE; Cutri et al. (2021).
@@
Table 4: Planetary physical parameters determined from our global MCMC analysis (medians and $1\sigma$ error bar) for TOI-5720 b, TOI-6008 b, and TOI-6086 b. a The equilibrium temperature derived assuming a null albedo. b Planetary mass and RV semi-amplitude values derived from RVs measurements collected by Subaru/IRD and TrES spectrographs for TOI-6008 b, and TOI-5720 b, respectively. c Predicted mass and RV semi-amplitude values calculated using the mass-radius relationship from Chen & Kipping (2017). d TSM (Transmission spectroscopy metric) and ESM (emission spectroscopy metric) values calculated from Kempton et al. (2018).
@@
Table 5: Derived stellar parameters ($\rho_{\star}$, $M_{\star}$, $R_{\star}$, $L_{\star}$, and $T_{\rm eff}$) and limb-darkening coefficients for TOI-5720, TOI-6008, and TOI-6086 based on our global analysis.
@@ 3.2 IRTF/SpeX spectroscopic analysis
3.3 Spectral energy distribution analysis and evolutionary models
We conducted an independent analysis of the broadband spectral energy distribution (SED) of the stars together with the Gaia DR3 parallax (with no systematic offset applied; see, e.g., Stassun & Torres 2021) to derive an empirical measurement of the stellar radii, following the same method as described in Stassun & Torres (2016); Stassun et al. (2017, 2018a). We pulled the 2MASS $JHK_{S}$ magnitudes, the WISE W1–W4 magnitudes, the Pan-STARRS $grizy$ magnitudes, and the Gaia $G_{\rm BP}G_{\rm RP}$ magnitudes. Together, the available photometry spans the full stellar SED over the wavelength range 0.4–20 $\mu$m (see Figure 8).
Figure 8: Spectral energy distribution (SED) fit of TOI-5720 (left), TOI-6008 (middle) and TOI-6086 (right). The gray curves are the best-fitting PHOENIX atmosphere models, the colored symbols with error bars are the observed fluxes (horizontal bars represent the effective bandpasses), and black symbols are the model fluxes.
We performed a fit using PHOENIX stellar atmosphere models (Husser et al. 2013a), with the free parameters being the metallicity ([Fe/H]) and effective temperature ($T_{\rm eff}$). We set null extinction due to the close proximity of the stars to Earth. The resulting fits (Figure 8) have a best-fit $T_{\rm eff}=3325\pm 75$ K, $T_{\rm eff}=3075\pm 75$ K, and $T_{\rm eff}=3200\pm 75$ K for TOI-5720, TOI-6008, and TOI-6086, respectively, and a best-fit [Fe/H] = $0.0\pm 0.3$, $+0.3\pm 0.3$, and $-0.2\pm 0.3$ respectively, with a reduced $\chi^{2}$ of 1.8, 1.9, and 2.4 respectively (the latter star exhibits some UV excess that may indicate chromospheric activity). Integrating the model SED gives the bolometric flux at Earth, $F_{\rm bol}=4.027\pm 0.094\times 10^{-10}$ erg s-1 cm-2, $F_{\rm bol}=2.839\pm 0.101\times 10^{-10}$ erg s-1 cm-2, and $F_{\rm bol}=2.051\pm 0.072\times 10^{-10}$ erg s-1 cm-2, respectively. Taking the $F_{\rm bol}$ together with the Gaia parallax directly gives the bolometric luminosity as $L_{\rm bol}=0.01615\pm 0.00019$ L, $L_{\rm bol}=0.004691\pm 0.000082$ L, and $L_{\rm bol}=0.00634\pm 0.00011$ L respectively. The stellar radii then follow from the Stefan-Boltzmann relation as $R_{\star}=0.383\pm 0.018$ R, $R_{\star}=0.242\pm 0.013$ R, and $R_{\star}=0.259\pm 0.013$ R, respectively. Finally, we can estimate the stellar mass from the empirical mass-luminosity relation of Mann et al. (2019), giving $M_{\star}=0.383\pm 0.019$ M, $M_{\star}=0.230\pm 0.011$ M, and $M_{\star}=0.254\pm 0.013$ M, respectively.
@@ 4.1 Data validation report for TESS
As we described in Section 2.1, the SPOC pipeline extracted the photometric measurements for TOI-5720, TOI-6008 and TOI-6086 using a cadence of 2-minutes, yielding orbital periods of $P=1.43446$ days, $P=0.85737$ day, and $P=1.38886$ days, and signal-to-noise ratios of $S/N=7.7$, $S/N=14.9$, and $S/N=8.1$, respectively. The TOI vetting team reviewed the SPOC data validation reports (Jenkins 2002; Jenkins et al. 2010, 2020) on May 26 2022, Sept 16 2022, and Sept 15 2022 (Guerrero et al. 2021) for TOI-5720, TOI-6008, and TOI-6086, respectively. The transit depths were found to be $1.03\pm 0.16$ ppt (for TOI-5720.01), $2.05\pm 0.27$ ppt (for TOI-6008.01), and $1.37\pm 0.24$ ppt (for TOI-6086.01), corresponding to a planet radius of $R_{p}=1.4\pm 0.3\pm$ R, $R_{p}=1.2\pm 0.2\pm$ R, and $1.0\pm 1.7\pm$ R for TOI-5720.01, TOI-6008.01, and TOI-6086.01, respectively. All additional validation tests, including bootstrap, centroid offset, ghost, difference-imaging centroid tests, and a search for discrepancies between odd and even transit depths were successfully passed. These conclusive tests prompted us to conduct further high-precision photometric follow-up and analysis to validate the planet candidates.
@@ 4.2 Vetting
While the transit method is one of the most fruitful technique for discovering exoplanets, several scenarios other than a transiting exoplanet can produce periodic dips in the brightness of a star. Numerous false-positive scenarios exist in astronomical observations: i) stellar binary systems, where two stars sharing a common center might create a signal resembling a planetary transit when their sizes differ or when their eclipse is partial; ii) blended binary stars, wherein the target blends the light of an eclipsing binary obscuring the secondary eclipse and confusing the primary with a planetary transit; and iii) other misleading signals from astrophysical sources (stellar spots, pulsations, and rotation) or instrument-related factors (jitter noise, and momentum dumps).
@@
Thus, the first essential step toward the confirmation of a transiting exoplanetary signal is the vetting. In addition to the SPOC DV reports for TOI-5720.01, TOI-6008.01, and TOI-6086.01, we conducted a uniform vetting analysis through the tool called discovery and vetting of exoplanets (DAVE; Kostov et al. 2019), which allowed us to examine a transit event on a double-level analysis: the pixel level by a photocenter analysis, and the light-curve level by a flux time-series analysis. The centroid module generates a difference image by subtracting the overall in-transit image from the out-of-transit image. It determines the photocenter by fitting the TESS pixel response function to the image, which is performed for each detected transit. The average centroid position of the events is calculated, along with its statistical significance. This aids in the vetting process. An offset in the centroid position from the expected target star position flags the event as a false positive (FP). However, the challenging interpretation due to artifacts or low S/N in DAVE-produced difference images makes the photocenter analysis unreliable. In these cases, we classify the signal as a planet candidate. The Modelshift module produces a phase-folded light curve with the best-fit trapezoid transit model. Its primary goal is to determine whether the signal source is an eclipsing binary system. It showcases key transit features, including the average primary signal and other significant features. Additionally, it compares odd and even transits by measuring their statistical difference. It also evaluates the transit shape, offering a comprehensive assessment of the observed transit properties.
@@
In the vetting analysis of planetary transit candidates, the reliability can be compromised when the S/N of the event is particularly low. This occurs because low S/N signals may be indistinguishable from noise or other sources of variability in the data, making it challenging to accurately identify genuine transit events. As a result, the vetting efforts may be less effective in confirming or ruling out potential candidates for these cases. TOI-5720.01, TOI-6008.01, and TOI-6086.01 are three low S/N planet candidates, with the first and the third being near the detection threshold (S/N=$7.1$, Borucki et al. 2011). However, their transit depths in the phase-folded light curve are slightly above the noise level. Moreover, we did not flag any particular anomalies in the shape of the transit, nor did statistical significant evidence arise between the odd and even transits. In the centroid module is concerned, TOI-5720.01 passed the photocenter test without any concern. The same test was inconclusive for both TOI-6008.01 and TOI-6086.01 since the TESS aperture masks used to extract their light curves are crowded with contaminating stars, resulting in an unreliable overall centroid measurement. However, for a few single clean centroid measurements the light photocenters are consistent with the position of both TOI-6008 and TOI-6086, as we discuss in Sect.4.3.